# UI-R1: Enhancing Action Prediction of GUI Agents by Reinforcement Learning

### v3

@@ 1 Introduction
Rule-based reinforcement learning or reinforcement fine-tuning (RFT) has recently emerged as an efficient and scalable alternative to SFT for the development of LLMs, which efficiently fine-tune the model with merely dozens to thousands of samples to excel at domain-specific tasks. It uses predefined task-specific reward functions, eliminating the need for costly human annotations. Recent works, such as DeepSeek-R1 (Guo et al., 2025), demonstrate the effectiveness of rule-based RL in mathematical problem solving by evaluating the correctness of the solution, while others (Liu et al., 2025b; Wang et al., 2025; Peng et al., 2025; Chen et al., 2025a; Huang et al., 2025; Zhou et al., 2025; Chen et al., 2025b) extend the algorithm to multimodal models, achieving notable improvements in vision-related tasks such as image grounding and object detection. By focusing on measurable objectives, rule-based RL enables practical and versatile model optimization across both textual and multimodal domains, offering significant advantages in terms of efficiency, scalability, and reduced reliance on large datasets.
@@
Regarding data preparation, we follow Muennighoff et al. (2025) and select just 130+ training mobile samples according to three criterion: difficulty, diversity, and quality, making our method remarkably data-efficient. Experiments demonstrate that UI-R1 achieves significant performance improvements on out-of-domain gounding tasks like ScreenSpot-Pro (Li et al., 2025) and computer scenarios in ScreenSpot (Cheng et al., 2024), indicating the potential of rule-based RL to tackle complex GUI-related tasks across diverse domains effectively.
@@ 2.2 Rule-Based Reinforcement Learning
Rule-based reinforcement learning has recently emerged as an efficient alternative to traditional training paradigms by leveraging predefined rule-based reward functions to guide model behavior. DeepSeek-R1 (Guo et al., 2025) first introduced this approach, using reward functions based on predefined criteria, such as checking whether an LLM’s final answer matches the ground truth for math problems. The reward focuses solely on the final results, leaving the reasoning process to be learned by the model itself.  Zeng et al. (2025) reproduces the algorithm on models with smaller sizes and illustrates its effectiveness on small language models. Subsequent works (Chen et al., 2025a; Shen et al., 2025; Liu et al., 2025b; Wang et al., 2025; Peng et al., 2025; Meng et al., 2025), extended the paradigm to multimodal models by designing task-specific rewards for visual tasks, including correct class predictions for image classification and IoU metrics for image grounding and detection. These studies demonstrate the adaptability of rule-based RL for both pure-language and multimodal models. By focusing on task-specific objectives without requiring extensive labeled datasets or human feedback, rule-based RL shows strong potential as a scalable and effective training paradigm across diverse tasks.
@@ 3.1 Preliminary
where $Mean$ and $Std$ represent the mean and standard deviation of the rewards, respectively. This normalization step ensures that responses are compared within the context of the group, allowing GRPO to better capture nuanced differences between candidates. Policy updates are further constrained by minimizing the KL divergence between the updated and reference models, ensuring stable RL learning.
@@ 3.2 Rule-Based Action Rewards
The rule-based reward function introduced by DeepSeek-R1 (Guo et al., 2025) represents a foundational step in rule-based RL by simply evaluating whether model predictions exactly match ground-truth answers. This straightforward approach efficiently aligns models with preference alignment algorithms and provides clear optimization signals. For vision-related tasks, works such as VLM-R1 (Shen et al., 2025) and Visual-RFT (Liu et al., 2025b) extend this idea by designing task-specific rewards. For image grounding tasks, they compute the IoU between the predicted and ground-truth bounding boxes as the reward. Similarly, for image classification tasks, rewards are determined by checking whether the predicted and ground-truth classes match.
@@ 4.1 GUI Grounding Capability
We assess the grounding capability of UI-R1 using two benchmarks: ScreenSpot (Cheng et al., 2024) and ScreenSpot-Pro (Li et al., 2025). ScreenSpot evaluates GUI grounding capability across mobile, desktop, and web platforms, while ScreenSpot-Pro focuses on high-resolution professional environments, featuring expert-annotated tasks spanning 23 applications, five industries, and three operating systems. Evaluation results of ScreenSpotV2 (Wu et al., 2024) are in Appendix B.
Model Method Model Size Data Size Web Desktop Average
Icon Text Icon Text
@@
SeeClick SFT 9.6B 1M 32.5 55.7 30.0 72.2 49.0
CogAgent SFT 18B - 28.6 70.4 20.0 74.2 51.0
Qwen2.5-VL SFT 3B 500 63.1 78.3 46.4 85.0 70.1
UGround-V1 SFT 7B 10M 70.4 80.4 63.6 82.5 75.2
AGUVIS SFT 7B 1M 70.7 88.1 74.8 85.7 80.4
@@
Qwen2-VL ZS 7B 0 25.7 35.2 54.3 76.3 46.5
Qwen2.5-VL ZS 3B 0 43.2 60.0 40.0 80.9 57.1
UI-R1 RFT 3B 136 73.3 85.2 59.3 90.2 78.6
@@ Appendix B Other Evaluation
B.1 ScreenSpot-V2
We also evaluate the model performance on ScreenSpot-V2 (Wu et al., 2024) and the results are in Table 6.
Model GUI specific Size Mobile Web Desktop Avg
Icon Text Icon Text Icon Text
SeeClick Yes 9.6B 50.7 78.4 32.5 55.2 29.3 70.1 55.5
OS-Atlas Yes 4B 59.7 87.2 63.1 85.9 46.4 72.7 71.9
OS-Atlas Yes 7B 75.8 95.2 77.3 90.6 63.6 90.7 84.1
UI-TARS Yes 2B 79.1 95.2 78.3 87.2 68.6 90.7 84.7
Qwen2.5-VL Framework
Qwen2.5-VL No 3B 66.8 92.1 46.8 72.6 44.3 83.0 70.4
Qwen2.5-VL No 7B 80.6 95.9 70.0 87.2 59.3 89.2 82.6
UI-R1(Ours) Yes 3B 84.3 96.2 75.4 89.2 63.6 92.3 85.4
B.2 Reasoning

### v4

# UI-R1: Enhancing Efficient Action Prediction of GUI Agents by Reinforcement Learning

@@ 1 Introduction
Rule-based reinforcement learning or reinforcement fine-tuning (RFT) has recently emerged as an efficient and scalable alternative to SFT for the development of LLMs, which efficiently fine-tune the model with merely dozens to thousands of samples to excel at domain-specific tasks. It uses predefined task-specific reward functions, eliminating the need for costly human annotations. Recent works, such as DeepSeek-R1 (Guo et al., 2025), demonstrate the effectiveness of rule-based RL in mathematical problem solving by evaluating the correctness of the solution, while others (Liu et al., 2025b; Wang et al., 2025b; Peng et al., 2025; Chen et al., 2025a; Huang et al., 2025; Zhou et al., 2025; Chen et al., 2025b) extend the algorithm to multimodal models, achieving notable improvements in vision-related tasks such as image grounding and object detection. By focusing on measurable objectives, rule-based RL enables practical and versatile model optimization across both textual and multimodal domains, offering significant advantages in terms of efficiency, scalability, and reduced reliance on large datasets.
@@
Regarding data preparation, we follow Muennighoff et al. (2025) and select just 130+ training mobile samples according to three criterion: difficulty, diversity, and quality, making our method remarkably data-efficient. Experiments demonstrate that UI-R1 achieves significant performance improvements on out-of-domain gounding tasks like ScreenSpot-Pro (Li et al., 2025a) and computer scenarios in ScreenSpot (Cheng et al., 2024), indicating the potential of rule-based RL to tackle complex GUI-related tasks across diverse domains effectively.
@@
  • We additionally develop an optimized version, UI-R1-E-3B, which significantly improves both grounding efficiency and accuracy.
@@ 2.2 Rule-Based Reinforcement Learning
Rule-based reinforcement learning has recently emerged as an efficient alternative to traditional training paradigms by leveraging predefined rule-based reward functions to guide model behavior. DeepSeek-R1 (Guo et al., 2025) first introduced this approach, using reward functions based on predefined criteria, such as checking whether an LLM’s final answer matches the ground truth for math problems. The reward focuses solely on the final results, leaving the reasoning process to be learned by the model itself.  Zeng et al. (2025) reproduces the algorithm on models with smaller sizes and illustrates its effectiveness on small language models. Subsequent works (Chen et al., 2025a; Shen et al., 2025a; Liu et al., 2025b; Wang et al., 2025b; Peng et al., 2025; Meng et al., 2025), extended the paradigm to multimodal models by designing task-specific rewards for visual tasks, including correct class predictions for image classification and IoU metrics for image grounding and detection. These studies demonstrate the adaptability of rule-based RL for both pure-language and multimodal models. By focusing on task-specific objectives without requiring extensive labeled datasets or human feedback, rule-based RL shows strong potential as a scalable and effective training paradigm across diverse tasks.
Efficient Reasoning.
Recent studies have sought to improve reasoning efficiency in large reasoning models (LRMs), addressing challenges such as redundant content (e.g., repeated definitions), over-analysis of simple problems, and shallow exploration of reasoning paths in complex tasks (Qu et al., 2025). To this end, several works (Xiao et al., 2025; Shen et al., 2025b; Aggarwal & Welleck, 2025; Li et al., 2025b; Wang et al., 2025a) introduce a length reward within reinforcement learning frameworks, which encourages concise and accurate reasoning by rewarding short, correct answers and penalizing lengthy or incorrect ones. Our work demonstrates that easy tasks like GUI grounding do not need reasoning process and UI-R1-E-3B achieves SOTA performance with fast grounding.
@@ 3.1 Preliminary
where $Mean$ and $Std$ represent the mean and standard deviation of the rewards, respectively. This normalization step ensures that responses are compared within the context of the group, allowing GRPO to better capture nuanced differences between candidates.
The policy model is then optimized by maximizing the following KL objective:
$\begin{split}\mathcal{J}_{GRPO}(\theta)=\mathbb{E}_{q\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}}(O|q)}\Bigg{[}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\Bigg{\{}\min\Bigg{[}\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{old}}(o_{i,t}|q,o_{i,<t})}\hat{A}_{i,t},\quad\\ \text{clip}\left(\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{old}}(o_{i,t}|q,o_{i,<t})},1-\epsilon,1+\epsilon\right)\hat{A}_{i,t}\Bigg{]}-\beta\mathbb{D}_{KL}\left[\pi_{\theta}||\pi_{ref}\right]\Bigg{\}}\Bigg{]}\end{split}$ (2)
where $\pi_{\theta}$ and $\pi_{old}$ are the current and old policy, and $\epsilon$ and $\beta$ are hyper-parameters introduced in PPO.
@@ 3.2 Rule-Based Action Rewards
The rule-based reward function introduced by DeepSeek-R1 (Guo et al., 2025) represents a foundational step in rule-based RL by simply evaluating whether model predictions exactly match ground-truth answers. This straightforward approach efficiently aligns models with preference alignment algorithms and provides clear optimization signals. For vision-related tasks, works such as VLM-R1 (Shen et al., 2025a) and Visual-RFT (Liu et al., 2025b) extend this idea by designing task-specific rewards. For image grounding tasks, they compute the IoU between the predicted and ground-truth bounding boxes as the reward. Similarly, for image classification tasks, rewards are determined by checking whether the predicted and ground-truth classes match.
@@
3.3 Fast Grounding
Our optimized grounding version, UI-R1-Efficient-3B, is trained in two stages: DAST training followed by NOTHINK training, with each stage lasting 4 epochs.
DAST
refines rule-based rewards by integrating the deviation between actual response length and the Token Length Budget (TLB) metric, allowing the reward to capture both task difficulty and length properties for effective difficulty-adaptive training (Shen et al., 2025b).
TLB metric is defined as:
$L_{budget}=p\cdot L_{\overline{r}}+(1-p)\cdot L_{max},p=\frac{c}{N}$ (5)
where $c$ is the number of correct responses sampled for a given question, and $N$ is the total number of sampled responses. $L_{\overline{r}}$ denotes the average token length of the correct responses, and $L_{\mathrm{max}}$ represents the maximum generation length.
The length reward $R_{\mathcal{L}}$ is defined as:
$R_{\mathcal{L}}=\begin{cases}\max(-0.5\lambda+0.5,\ 0.1)&\text{if correct}\\ \min(0.9\lambda-0.1,\ -0.1)&\text{if incorrect}\end{cases}\quad\text{where}\quad\lambda=\dfrac{L_{i}-L_{budget}}{L_{budget}}$ (6)
The reward in Equation 3 is then calibrated by the length reward:
$R=R_{\mathcal{T}}+R_{\mathcal{C}}+R_{\mathcal{F}}+{\color[rgb]{1,0.75390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.75390625,0}\textit{$R_{\mathcal{L}}$}},$ (7)
NOTHINK
enables fast and direct grounding by removing the $<$$\boldsymbol{answer}$$>$ tags, thereby bypassing explicit reasoning steps during both training and inference (Li et al., 2025b).
The corresponding ablation results are presented in Table 5. Based on these results, we can conclude that
”Reasoning is not essential for simpler tasks, such as GUI grounding.”
@@ 4.1 GUI Grounding Capability
We assess the grounding capability of UI-R1 using two benchmarks: ScreenSpot (Cheng et al., 2024) and ScreenSpot-Pro (Li et al., 2025a). ScreenSpot evaluates GUI grounding capability across mobile, desktop, and web platforms, while ScreenSpot-Pro focuses on high-resolution professional environments, featuring expert-annotated tasks spanning 23 applications, five industries, and three operating systems. We also evaluate the model performance on ScreenSpot-V2 (Wu et al., 2024) and the results are in Table 2.
Model Method Model Data Mobile Web Desktop Average
Size Size Icon Text Icon Text Icon Text
@@
CogAgent SFT 18B - 24.0 67.0 28.6 70.4 20.0 74.2 47.4
SeeClick SFT 9.6B 1M 52.0 78.0 32.5 55.7 30.0 72.2 53.4
UGround-V1 SFT 7B 10M 60.3 82.8 70.4 80.4 63.6 82.5 73.3
Qwen2.5-VL SFT 3B 500 71.2 95.2 63.1 78.3 46.4 85.0 75.7
AGUVIS SFT 7B 1M 78.2 88.3 70.7 88.1 74.8 85.7 81.8
@@
Qwen2-VL ZS 7B 0 60.7 75.5 25.7 35.2 54.3 76.3 55.3
Qwen2.5-VL ZS 3B 0 61.1 90.5 43.2 60.0 40.0 80.9 65.0
UI-R1-3B (Ours) RFT 3B 136 84.7 95.6 73.3 85.2 59.3 90.2 83.3
UI-R1-E-3B (Ours) RFT 3B 2K 83.0 97.1 85.0 91.7 77.9 95.4 89.2
@@
Model GUI specific Size Mobile Web Desktop Avg
Icon Text Icon Text Icon Text
SeeClick Yes 9.6B 50.7 78.4 32.5 55.2 29.3 70.1 55.5
OS-Atlas Yes 4B 59.7 87.2 63.1 85.9 46.4 72.7 71.9
OS-Atlas Yes 7B 75.8 95.2 77.3 90.6 63.6 90.7 84.1
UI-TARS Yes 2B 79.1 95.2 78.3 87.2 68.6 90.7 84.7
Qwen2.5-VL Framework
Qwen2.5-VL No 3B 66.8 92.1 46.8 72.6 44.3 83.0 70.4
Qwen2.5-VL No 7B 80.6 95.9 70.0 87.2 59.3 89.2 82.6
UI-R1-3B (Ours) Yes 3B 84.3 96.2 75.4 89.2 63.6 92.3 85.4
UI-R1-E-3B (Ours) Yes 3B 83.9 98.2 83.7 93.2 75.0 94.8 89.5
@@ 4.4 Ablation Study
Fast Grounding
The ”First DAST, then NOTHINK” approach enables the model to learn grounding by gradually transitioning from slower, more thoughtful reasoning to faster, more direct, and accurate responses. The order of these steps is crucial, as disrupting their sequence or removing one of them may significantly impact training efficiency. Moving forward, we aim to explore whether reasoning, as embodied in the thinking phase, is essential for handling high-level planning tasks, or if simpler, direct approaches can yield similar performance in these more complex contexts.
[…]
DAST + NOTHINK (UI-R1-E-3B) 89.5 33.5 0
NOTHINK + DAST 87.7 31.6 0
DAST 87.3 30.8 short
NOTHINK 86.4 32.1 0
THINK V2 85.7 29.8 long
THINK V1 (UI-R1-3B) 85.4 17.8 long
Table 5: Abalation of thinking-or-not training methods (DAST, NOTHINK or THINK) on grounding tasks ScreeSpot V2 (SSV2) and ScreenSpot Pro (SSPro). The total training epoches are all set as 8.
@@ Appendix B Other Evaluation
B.1 Reasoning